These were accepted.
A draft has been circulated to CEG
Services-unit are moving this, but their timetable may not fit with
the need to get the atabeast off the FibreChannel fabric and directly
central. One option might be to access the
remaining data from
central after the connection arrangements
A meeting was imminent, though there hadn't been any papers circulated yet. Alastair will raise both our items at the meeting. (Actions left, pending a followup report.)
This is almost done.
Responses now need to be correlated and non-responders sent a reminder. Ultimately it's up to HoS to ensure enforcement -- we'll report in some to-be-determined-later way.
The wiki page is tracking activity. Progress has been reasonably steady. As above, we'll report to HoS on what's what.
Alastair discovered that IS's response to his email had vanished into UniDesk. Having found it, he reported that they have a membership for their UNIX group only. He'll now set up a separate membership for us, on a trial basis. Action amended.
This is now mostly done. Our queues are not visible to all and sundry any more.
Actions added (see below for details):
centralthat haven't been done recently.
This is being progressed as part of the SL6 upgrade programme.
Blog articles completed:
Blog articles suggested:
Report from Computing Executive Group
Reports from units.
The Services and Infrastructure Units would both like to have a presence while the cleaning is done. George will respond to Alan, copying in Craig and Ian. The holiday season is almost upon us, so it would be better not to delay.
(Followup remarks: Alan has now scheduled this for 18th June at 10:00. He has also sent a copy of the contractor's quote and method statement. This is not suitable for open circulation, but can be passed to those who need to know.)
Toby and Neil will arrange for the new certificates to be installed, and for any from the "additional" list to be obtained within the non-commercial-use constraints.
Chris asked that anyone who has a VM on
shut it down and restart it if they haven't done so recently.
To clarify, it's not currently the plan to replace flipdesk machines with new ones, as they're generally less used than the main lab machines. It's more likely that they'll be replaced with machines displaced from elsewhere.
Topics for discussion
This is being tracked by the actions previously raised.
/public-- essentially any automounted NFS space.
rfe-driven side of the map generation process was working as it had always done, something in the move to Prometheus had broken the part which takes the posixAccount data as its source. It still runs successfully by hand, though, when it is realised that such is required.
/homemap. After some discussion it was concluded that we should keep this, as although it's not well-publicised it is mentioned in a few places and there are likely to be long-standing scripts which rely on it.
autofswas discussed only in passing here, though it has been mooted in the past.
ldapBuildAmdMaps. He'll take a look.
Alison gave a short presentation on the recent student survey, which produced some interesting results. Some discussion ensued. There's a summary report here, and the individual "is there anything else" reponses are here.
The next meeting would normally be on Wednesday 13th June 2012, but watch this space...
Please contact us with any comments or corrections.
Unless explicitly stated otherwise, all material is copyright The University of Edinburgh